ATTEMPT ALL THE QUESTIONS

(4 ) As cost accountant of Rackets Ltd, you estimate the standard cost of making one unit is as follows:

Direct material 4 kilos at £7 per kilo

Direct wages 5 hours at £10 per hour

The actual cost of a batch of 1000 units was:

Direct material £28960 (4328 kilos)

Direct wages £50200 (5370 hours)

Calculate the following:

(vi) The total labour cost variance

(vii) The total cost variance
Answer #1

Given data,

Estimated standard cost of making one unit

Direct material : 7 per kilo

Direct wages : 10 per hour

Actual cost of making 1000 unit:

Actual Direct Wages : 50200

Solution:

Standard Cost of making 1000 unit :

Standard Direct Material : 1000*4*7 = 28000

Standard Direct Wages : 1000*5*10 = 50000

(Vi) Total Labour Cost Variance = (Standard Hour * Standard Rate ) - (Actual Hour * Actual Rate)

= Standard Direct Wages - Actual Direct Wages

= 50000 - 50200 = 200 (Adverse)

(vii) Total Cost Variance = Total Material Cost Variance + Total Labour Cost Variance

(a) Total Material Cost Variance = (Standard Unit * Standard Rate) - (Actual Unit * Actual Rate)

= Standard Material Cost - Actual Material Cost

= 28000 - 28960 = 960 (Adverse)

(b) Total Labour Cost Variance (as per solution no. (vi) above = 200 (Adverce)

Total Cost Varialnce = Total Material Cost Variance + Total Labour Cost Variance

= 200+960 = 1160 (Adverse)